OpenStudy (anonymous):

Write the series in expanded form and find the sum:

OpenStudy (anonymous):

\[\sum_{J=1}^{5}(4j-16)\]

OpenStudy (ikram002p):

so can u find the sum nw ?

OpenStudy (neer2890):

for 1st term put 1 instead of j so 1st term is(4*1-16)=4-16=-12 for second term put 2 instead of j and find u

OpenStudy (neer2890):

find all terms upto j=1 to j=5

OpenStudy (anonymous):

(4 ∙ 1 - 16) (4 ∙ 2 - 16) (4 ∙ 3 - 16) (4 ∙ 4 - 16) (4 ∙ 5 - 16) = -20 (4 ∙ 1 - 16) + (4 ∙ 2 - 16) + (4 ∙ 3 - 16) + (4 ∙ 4 - 16) + (4 ∙ 5 - 16) = -20 (4 ∙ 1 - 16) + (4 ∙ 5 - 16) = -8 (4 ∙ 1 - 16) (4 ∙ 5 - 16) = -16

OpenStudy (neer2890):

you don't have to multiply all\[\sum_{1}^{5}\] is for sum of terms from 1 to 5

OpenStudy (ikram002p):

there is so much easy way \(\sum_{j=1}^{5}(4j-16)=4\sum_{j=1}^{5}(j)-\sum_{j=1}^{5}( - 16)\)

OpenStudy (neer2890):

@ikram002p is correct but i think it should be + instead of - between two summations.

OpenStudy (ikram002p):

sry typo :D \(\sum_{j=1}^{5}(4j-16)=4\sum_{j=1}^{5}(j)-\sum_{j=1}^{5}(16)\)

OpenStudy (anonymous):

So which is it? A, B, C, or D?

OpenStudy (anonymous):

because none of what you have typed so far matches with any of the answers that I posted....

OpenStudy (neer2890):

just plug in the values of j and find out yourself.

OpenStudy (ikram002p):

ok lets try to know which one is wrong for A (4 ∙ 1 - 16) (4 ∙ 2 - 16) (4 ∙ 3 - 16) (4 ∙ 4 - 16) (4 ∙ 5 - 16) = -20 the operation btw term is * it should be + so not A

OpenStudy (neer2890):

to find the correct answer of your options, you have to adopt the long method which i have told you earlier.

OpenStudy (ikram002p):

D is wrong as well , since it took first+last term and * operation two

OpenStudy (anonymous):

So that eliminates (4 ∙ 1 - 16) (4 ∙ 5 - 16) = -16 too right?

OpenStudy (ikram002p):

yep ur right :) nw lets see C

OpenStudy (ikram002p):

(4 ∙ 1 - 16) + (4 ∙ 5 - 16) = -8 it tool the first term and the last term sum but we have from 1 to 5 means 5 terms so not right

OpenStudy (anonymous):

So it's B (4 ∙ 1 - 16) + (4 ∙ 2 - 16) + (4 ∙ 3 - 16) + (4 ∙ 4 - 16) + (4 ∙ 5 - 16) = -20

OpenStudy (ikram002p):

so nw u have B , but make sure they did the correct sum (4 ∙ 1 - 16) + (4 ∙ 2 - 16) + (4 ∙ 3 - 16) + (4 ∙ 4 - 16) + (4 ∙ 5 - 16) = -20

OpenStudy (anonymous):

Thank you so much @ikram002p! You made it very simple and easy for me to understand!

OpenStudy (ikram002p):

(4 ∙ 1 - 16) =-12 (4 ∙ 2 - 16)= -8 (4 ∙ 3 - 16) =-4 (4 ∙ 4 - 16) =0 (4 ∙ 5 - 16) = 4 the sum =-12+-8+-4+-0+4=-20

OpenStudy (ikram002p):

np :)

OpenStudy (vishweshshrimali5):

\(\color{blue}{\text{Originally Posted by}}\) @ikram002p sry typo :D \(\sum_{j=1}^{5}(4j-16)=4\sum_{j=1}^{5}(j)-\sum_{j=1}^{5}(16)\) \(\color{blue}{\text{End of Quote}}\) Well let me give you an alternative and very short method. I am going to continue from this step of @ikram002p : \(\large{\sum_{j=1}^{5}(4j-16)=4\sum_{j=1}^{5}(j)-\sum_{j=1}^{5}(16)}\)

OpenStudy (vishweshshrimali5):

See: observe the first summation: \[\large{\sum_{j = 1}^{5}(j)}\] It is like adding first 5 natural numbers. Now, I am providing you a formula: sum of first n natural numbers = \(\cfrac{n(n+1)}{2}\) Its very easy to remember.

OpenStudy (vishweshshrimali5):

So, sum of first 5 natural numbers = \(\cfrac{5(5+1)}{2}\) Can you solve this and tell me the value ?

OpenStudy (vishweshshrimali5):

@sapphirelearner1995 ?

OpenStudy (anonymous):

\[30/2=15\]

OpenStudy (vishweshshrimali5):

Good. Now, replace this value by first summation. Now, for the second summation: \(\large{\sum_{j = 1}^{5} (16)}\) Here, 16 is a constant and does not depend on j. So, I can take it outside and I will be left with: \(\large16{\sum_{j = 1}^{5} (1)}\) Now, this is equivalent to first adding 1, 5 times and then multiplying the answer by 16. So, can you find this summation ?